Calculate the energy separations in joules, kilojoules per mole, and reciprocal centimeters between the levels (a) n = 2 and n = 1, (b) n = 6 and n = 5 of an electron in a box of length 2.0 nm
Answer #1

The energy of a particle in a box is given as follows: 8ml Here, n is the energy state, h is the Plancks constant (6.626x 10Therefore, the energy separation in kilojoule per mole is 27.22 kJ mol-. According to Plancks hypothesis, energy is given as(b) The general formula for energy separation is as follows: (nz-n?)h2 8m2 Here, n, is 6 and n, is 5. Putting all the valuesThe energy separation in kilojoule per moles is given as follows: 11 =10-kJ Therefore, X10-3 1-1 1.66x 10-19 1 - 1.66 x10-19Putting the values as follows: *)(3x109 m ) 1.66x10-19 * =(6.626x10-24 v=0.084x10 m =0.084x10 cm Therefore, the energy separa